If there are 127 full time employees and 17% absenteeism. And the work 5,800 inquiries per day and each person releases 5.2 inquire per hr per person how many written inquires can be released as a whole each day?
Answer #1

So there are 127 x (1-0, 17) = 127 x 0.83 persons present on each day on an average.

Assuming 8 hours a day, the number of enquiries can be released = 127 \times 0.83 \times 5.2 \times 8

\approx 4385

If you have doubst. ;let me know below. Thanks